OpenStudy (rebeccapink6398):

PLZ HELP HCECK MY WORK!!!!!!!!!!! 30% of a number is 120. What is 75% of the number? A. 90 B. 300 C. 48 D. 27 >>I think its B?

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

It's the only answer that makes sense. If 30% of the number is 120, then the number cannot by under 120.

OpenStudy (rebeccapink6398):

how about this one??>>>> Kirk's family drove 36 miles in 50 minutes, which was 48% of the distance to his grandmother's house. What was the total distance to Kirk's grandmother's house? A. 48 miles B. 24 miles C. 17.28 miles D. 75 miles think its A!!

OpenStudy (mathstudent55):

Use a proportion: 30% is to 120 as 75% is to x \( \dfrac{30}{120} = \dfrac{75}{x} \) 30x = 75 * 120 30x = 9000 x = 300 Answer is B. You are correct.

OpenStudy (mathstudent55):

36 miles is 48% x miles is 100% Use a proportion: \( \dfrac{36}{48} = \dfrac{x}{100} \) 48x = 36 * 100 48x = 3600 x = 75

OpenStudy (anonymous):

He is right, I read the question wrong.

OpenStudy (mathstudent55):

Another way of doing the second problem: 36 is 48% of x 36 = 48% * x 36 = 48/100 * x 36 = 48x/100 48x/100 = 36 48x = 3600 x = 75

OpenStudy (mathmale):

@RebeccaPink6398 : Would you please post each question separately. Also, please show what YOU have done before asking for help with a given problem.

OpenStudy (anonymous):

yes, the answer is B. the proportion for solving any problem like this is setting up a proportion. the rule is % over the overall number. So plug in what you know and name the missing piece x and solve. so the one for this problem would be 30 over 120 = 75 over x